# Are you doing better than random guessing? A call for using negative controls when evaluating causal discovery algorithms

### v1

@@ Abstract
New proposals for causal discovery algorithms are typically evaluated using simulations and a few select real data examples with known data generating mechanisms. However, there does not exist a general guideline for how such evaluation studies should be designed, and therefore, comparing results across different studies can be difficult. In this article, we propose a common evaluation baseline by posing the question: Are we doing better than random guessing? For the task of graph skeleton estimation, we derive exact distributional results under random guessing for the expected behavior of a range of typical causal discovery evaluation metrics (including precision and recall). We show that these metrics can achieve very large values under random guessing in certain scenarios, and hence warn against using them without also reporting negative control results, i.e., performance under random guessing. We also propose an exact test of overall skeleton fit, and showcase its use on a real data application. Finally, we propose a general pipeline for using random controls beyond the skeleton estimation task, and apply it both in a simulated example and a real data application.
@@ 1 Introduction
Causal discovery algorithms seek to infer information about a causal data generating mechanism by analyzing empirical data it generated. The causal data generating mechanism is typically represented by a causal graph, for example an equivalence class of directed acyclic graphs (DAGs). A highly productive research community has published a plethora of new causal discovery algorithms within the last 30 years or so. Naturally, this fast growing battery of available algorithms requires some standards and guidelines for evaluating and benchmarking their performance. Because the result of a causal discovery algorithm is an estimated graph (or family of graphs), rather than one or more scalars, it is not entirely straight-forward to use classic approaches for performance evaluation from neither machine learning nor statistics.
@@
Nonetheless, machine learning classification metrics originally developed for evaluating prediction tasks, are often used to evaluate causal discovery algorithms. Most commonly, precision and recall, or possibly their harmonic mean, the F1 score, are reported, although some studies also focus on other metrics, e.g. negative predictive value (Petersen et al., 2023b). These metrics are computed from graph-level confusion matrices summarizing either agreement on placement of oriented edges (primarily used for DAG discovery evaluation), adjacencies (i.e., edge placement without considering orientation), and/or arrowheads among correctly placed adjacencies (conditional orientation). Typically, they are reported as averages over numerous simulations. Sometimes the results are stratified by graphical parameters (e.g., true graph density), data-related parameters (e.g., sample size), or simply reported as averages across several such settings.
@@
Alternative metrics developed specifically for graphs also exist, where the structural hamming distance is the most widely used example, probably due to its cheap computation and easy interpretation. An alternative metric focusing more on the causal implications of the graphs is the structural intervention distance (SID) (Peters and Bühlmann, 2015), although it is most naturally suited for DAG-DAG comparison, and hence not readily applicable for all discovery evaluation tasks. A more recent proposal is the adjustment identification distance (Henckel et al., 2024), which also focuses on differences in causal inference based on the graph.
There are thus many different possible choices of metrics for evaluating causal discovery algorithms. However, no general guidelines exist for how to then interpret the values these metrics take. What is a high or low number? Of course, in a given evaluation study, multiple algorithms can be compared according to their relative performance. But it is not straight-forward to assess what values of the metrics should then be expected on just slightly different evaluation settings. This makes it very difficult to compare results across different evaluation studies with just marginally different designs.
In other sciences, e.g. epidemiology, negative controls are used to ensure the validity of a given study design (Lipsitch et al., 2010). A negative control is an experimental setting where the experimenters expect to see no effect. For example, we expect smoking to affect death by lung cancer, but probably not death by traffic accidents, and hence if we see an effect of smoking on death by traffic accidents in our study, there might be issues with the design. By analyzing data from the negative control scenario, and comparing it with the intended experiment, it may thus be possible to quantify unaccounted biases that are confounded by the experimental design. It is hence a pragmatic approach that acknowledges that we cannot always understand all aspects of a given scientific study design, and what may bias it, but we can instead sometimes specify what should not be able to affect our outcome of interest.
@@
We propose to use a negative control concept to obtain an interpretable benchmark for any causal discovery evaluation study: Namely, to investigate what values of the metrics of interest can be obtained using random guessing (a negative control). We discuss random controls in two different settings: First, we consider the task of skeleton estimation, i.e., estimating e.g. a DAG without any orientation information. In this case, we derive exact distributional results for the expected behavior under random guessing (Section 3), and we use these results to compute expected negative control values for a range of often-used metrics (Section 4). We furthermore propose an exact test of overall skeleton fit (Section 5, and provide an example of its use of real data. Secondly, we consider more general metrics, that are not just concerned with skeleton estimation, and propose a negative control pipeline for this case (Section 6). We provide two examples of its use, both in a simulation study and in a real data application.
@@
But before we turn to these general results, we provide an example of a case where well-known metrics such as adjacency precision and recall do perhaps not behave exactly as one would expect.
@@ 2 Precision and recall: A cautionary tale
Consider the two DAGs in Figure 1. The graph on the left (a) is the true DAG, and the graph on the right (b) is an estimated graph produced by a causal discovery procedure. We compute their adjacency confusion matrix in order to evaluate the performance of the discovery procedure:
@@
Estimate Adjacency $tp=6$ $fp=1$
Non-adjacency $fn=2$ $tn=1$
which results in
[…]
Are these numbers high or low? Although these values are not too far off from the performance of well-established causal discovery algorithms on simulated data (and much better than typical performance on real data), we will argue that they are indeed as low as can be for this specific discovery task, since the applied ”discovery algorithm” was simply random guessing. The DAG in Figure 1 (b) was drawn randomly among DAGs with 7 edges over 5 nodes (Erdös-Rényi type graph), and hence had absolutely no information about the true data generating mechanism.
The example provided here is not just the result of a conveniently chosen random seed. The example graphs were in fact chosen specifically to match the most typical performance when using a randomly drawn 5 node DAG with 7 edges to estimate the skeleton of a ”true” 5 node DAG with 8 edges. This can for example be demonstrated in a simulation study: We simulated 1000 random Erdös-Rényi type DAGs over 5 nodes each with 7 edges and used these ”random guesses” as estimates of the DAG in Figure 1 (a). This resulted in a median precision of 0.86 and a median recall of 0.75, i.e., numbers that exactly match the performance of the example just described.
Is it then a curious artefact for very dense graphs? Or ”small” graphs over e.g. 5 nodes? This is not the case. As we will show in the following section, the phenomenon does not depend on the number of nodes, and depending on the choice of metrics, can occur also in modestly dense graphs.
@@ 3 Distributional results for adjacency metrics under random guessing
Estimate Adjacency $TP$ $FP$ $m_{\text{est}}$
Non-adjacency $FN$ $TN$
@@
Table 1: A generically labelled adjacency confusion matrix. Entries marked with dashes are sums that will not be used for the derivations here.
Consider a DAG $G$ over $d$ nodes, and let $m_{\text{true}}$ denote the number of edges in $G$. Let $\hat{G}$ be another DAG over $d$ nodes used as an estimate of $G$, and let $m_{\text{est}}$ be the number of edges in $\hat{G}$. Finally, let $m_{\text{max}}=\sum_{i=1}^{d-1}i=\frac{1}{2}(d-1)d$ denote the maximal number of possible edges in a DAG over $d$ nodes (corresponding to a fully connected graph).
@@
We can describe the performance of $\hat{G}$ as an adjacency/skeleton estimator of $G$ through a (generic) confusion matrix as seen in Table 1. Note that for a causal discovery problem, $m_{\text{max}}$ and $m_{\text{true}}$ will generally be given a priori, as the number of nodes does not change, and the ground truth is considered fixed. Moreover, for many causal discovery procedures, it further makes sense to consider $m_{\text{est}}$ fixed – at least for a specific value of a tuning parameter (e.g. significance level for testing or penalty for a score) – as we do not often try to estimate the correct number of edges from data (see Sections 6 and 7 for considerations in cases where this latter assumption is not meaningful).
We now make the following important observation: If edges are placed uniformly in both $G$ and $\hat{G}$ (corresponding to an Erdös-Rényi type graph), edge placement can be seen as a random urn experiment with two colors of balls (say, blue and white), where white balls correspond adjacencies included in $G$, and blue balls are adjacencies not in $G$. A random causal discovery procedure will then metaphorically draw ”balls” (i.e., edges) randomly without replacement, and some will be correct (white), while others will be false (blue). Since the number of white balls ($m_{\text{true}}$), the number of draws ($m_{\text{est}}$) and the total number of balls ($m_{\text{max}}$) are all known a priori, the number of drawn white balls (true positive adjacencies) will by definition follow a hypergeometric distribution parameterized by $m_{\text{max}}$, $m_{\text{true}}$ and $m_{\text{est}}$:
$\text{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$
Note that this is an exact distributional result, not an asymptotic statement.
@@
This observation gives rise to several useful applications: First, we can compute the expected value, median and uncertainty estimates (e.g. confidence interval) for the number of true positive adjacencies under random guessing. Secondly, since we are also conditioning on $m_{\text{max}}$, $m_{\text{true}}$ and $m_{\text{est}}$, we can further compute expectations and draw statistical inference for any function of the confusion matrix, including precision, recall and F1. We provide formulas for these below in Section 4. Thirdly, we can construct an exact statistical test of overall skeleton fit by considering how much the number of true positives in a given estimated graph diverts from its expected distribution under a null hypothesis of random edge placement. We propose such a test in Section 5 below.
@@ 4 Expectations and quantiles of adjacency metrics under random guessing
Precision $\frac{m_{\text{true}}}{m_{\text{max}}}$ $\frac{k_{q}}{m_{\text{est}}}$
Recall $\frac{m_{\text{est}}}{m_{\text{max}}}$ $\frac{k_{q}}{m_{\text{true}}}$
F1 $\frac{2\cdot m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}\cdot m_{\text{est}}+m_{\text{max}}\cdot m_{\text{true}}}$ $\frac{2\cdot k_{q}}{m_{\text{est}}+m_{\text{true}}}$
NPV $1-\frac{m_{\text{true}}}{m_{\text{max}}}$ $\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+k_{q}}{m_{\text{max}}-m_{\text{est}}}$
Specificity $1-\frac{m_{\text{est}}}{m_{\text{max}}}$ $\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+k_{q}}{m_{\text{max}}-m_{\text{true}}}$
Table 2: Expected values and quantile expressions under random guessing for five commonly used metrics computed for adjacency performance. $q_{k}$ denotes the $k$th quantile from $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$.
Since $TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, by definition we have that
$\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}$
and by considering the quantile function of $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, we can find the expected median, or construct a confidence interval as e.g. the central 95% of the distribution.
Table 2 provides an overview of expected values and quantiles under random guessing for five metrics commonly used for evaluating adjacency placement for causal discovery algorithms, namely precision, recall, F1 score, negative predictive value (NPV) and specificity. As an example, we here showcase derivations for precision, and refer to Appendix A for derivations for the remaining four metrics.
@@
We first express precision as a function of $TP,m_{\text{max}},m_{\text{true}},m_{\text{est}}$:
$\text{precision}\operatorname{|}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{TP}{TP+FP}$ (1)
$=\frac{TP}{TP+m_{\text{est}}-TP}$ (2)
$=\frac{TP}{m_{\text{est}}}$ (3)
@@
$\text{E}(\text{precision}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{1}{m_{\text{est}}}\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
@@
The linearity in equation (3) also makes it easy to obtain e.g. a 95% exact confidence interval under the null hypothesis of random guessing, by simply applying the transformation from (3) to the appropriate quantiles of $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$. For example, an exact 95% confidence interval for precision under the null is given by
$\text{CI}_{0.95}=\left(\frac{1}{m_{\text{est}}}q_{(0.025,m_{\text{max}},m_{\text{true}},m_{\text{est}})},\frac{1}{m_{\text{est}}}q_{(0.975,m_{\text{max}},m_{\text{true}},m_{\text{est}})}\right)$
@@
$\text{median}(\text{precision}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{1}{m_{\text{est}}}q_{(0.5,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$
@@
Moreover, we note that under random guessing, the expected precision does not depend on the number of edges in the estimated graph ($m_{\text{est}}$), only on the number of edges in the true graph ($m_{\text{true}}$) and the maximal possible number of edges ($m_{\text{max}}$). But recall increases linearly as a function of the number of estimated edges. Hence, if we are using random guessing, a ”free lunch” in optimizing precision and recall can be had by simply estimating a very large number of edges, even including the trivial fully connected graph. This can also be seen from the expected value of the F1 score under random guessing, which increases monotonically with the number of estimated edges: It is always better to just add another edge.
@@
4.1 Example: Expected precision and recall for a dense 5 node DAG skeleton

### v2

@@ Abstract
New proposals for causal discovery algorithms are typically evaluated using simulations and a few selected real data examples with known data generating mechanisms. However, there does not exist a general guideline for how such evaluation studies should be designed, and therefore, comparing results across different studies can be difficult. In this article, we propose to use negative controls as a common evaluation baseline by posing the question: Are we doing better than random guessing? For the task of graph skeleton estimation, we derive exact distributional results under random guessing for the expected behavior of a range of typical causal discovery evaluation metrics, including precision and recall. We show that these metrics can achieve very favorable values under random guessing in certain scenarios, and hence warn against using them without also reporting negative control results, i.e., performance under random guessing. We also propose an exact test of overall skeleton fit, and showcase its use on a real data application. Finally, we propose a general pipeline for using negative controls beyond the skeleton estimation task, and apply it both in a simulated example and a real data application.
@@ 1 Introduction
Causal discovery algorithms seek to infer information about a causal data generating mechanism by analyzing empirical data it generated. The causal data generating mechanism is typically represented by a causal graph, for example an equivalence class of directed acyclic graphs (DAGs). A highly productive research community has published a plethora of new causal discovery algorithms within the last 30 years or so. Naturally, this fast growing battery of available algorithms requires some standards and guidelines for evaluating and benchmarking their performance. Because the result of a causal discovery algorithm is an estimated graph (or family of graphs), rather than one or more scalars, it is not entirely obvious how to use classic approaches for performance evaluation from neither machine learning nor statistics.
@@
Nonetheless, machine learning classification metrics originally developed for evaluating prediction tasks are often used to evaluate causal discovery algorithms. Most commonly, precision and recall, or possibly their harmonic mean, the F1 score, are reported, although some studies also focus on other metrics, e.g., negative predictive value [Petersen et al., 2023b]. These metrics are computed from graph-level confusion matrices summarizing either agreement on placement of oriented edges (primarily used for DAG discovery evaluation), adjacencies (i.e., edge placement without considering orientation), and/or arrowheads among correctly placed adjacencies (conditional orientation). Typically, they are reported as averages over numerous simulations. Sometimes the results are stratified by graphical parameters (e.g., true graph density), data-related parameters (e.g., sample size), or simply reported as averages across several such settings.
@@
Alternative metrics developed specifically for graphs also exist; the structural Hamming distance [Tsamardinos et al., 2006] is the most widely used example, probably due to its cheap computation and easy interpretation. A different metric focusing more on the causal implications of the graphs is the structural intervention distance (SID) [Peters and Bühlmann, 2015], although it is most naturally suited for DAG-DAG comparison, and hence not readily applicable for all discovery evaluation tasks. A more recent proposal is the adjustment identification distance [Henckel et al., 2024], which also focuses on differences in causal inference based on the graph, or the separation distance [Wahl and Runge, 2025], which counts agreement in separation statements. There are thus many different possible choices of metrics for evaluating causal discovery algorithms. However, no general guidelines exist on how to then interpret the values these metrics take: What is a high or low number?
An often-used strategy for answering this question in experimental sciences is to conduct a controlled experiment. In such an experiment, the intervention of interest (here: a causal discovery algorithm) is compared to a control condition. When this control condition cannot have any influence on the outcome of interest, it is denoted a negative control. For example, say we want to study the impact of a fertilizer on plant growth. We plant 100 seeds in two plots with similar conditions, except that one (the treatment group) receives the fertilizer, while the other (the negative control) does not. After, say, 10 days, we measure the heights of the plants, and we use the average difference in heights as a measure of the effect of the fertilizer. By including the negative control we obtain a direct measure of the specific effect of the treatment.
Alternatively, we could have compared the fertilizer of interest with another active treatment, perhaps an alternative well-known fertilizer (denoted a positive control). This comparison gives less information about the specific treatment of interest. For example, we cannot falsify a hypothesis saying that neither fertilizer has any effect. In some scientific fields, for example human drug trials, using positive controls is the only viable option, as it would be utterly unethical to deny patients treatment (if one exists) in order to obtain a negative control group for evaluating a new proposed treatment. But of course, no such concerns are relevant for evaluating causal discovery algorithms. Nonetheless, the current standard practice is to report positively controlled experiments: A new candidate algorithm is typically compared to a selection of existing algorithms. Such a comparison in itself does not provide information about whether either of the considered algorithms work. Moreover, because causal discovery evaluations are very sensitive towards experimental settings concerning graph sparsity (as we will demonstrate below), it is not straight-forward to generalize findings from such a positively controlled experiment to infer what performance should be expected on just slightly different evaluation settings. This makes it very difficult to compare results across different evaluation studies with just marginally different designs.
@@
We propose to use negative controls to obtain an interpretable benchmark for any causal discovery evaluation study: Namely, to investigate what values of the metrics of interest can be obtained using random guessing (a negative control), and report this alongside findings from positive controls (alternative algorithms). Others have reported results from a single random guess alongside causal discovery benchmarks [Lachapelle et al., 2020], but as we will argue, a more structured approach to including negative controls provides many benefits. We discuss negative controls in two different settings: First, we consider the task of skeleton estimation, that is estimating e.g., a DAG without taking orientation information into account. For this case, we derive exact distributional results for the expected behavior under random guessing (Section 3), and we use these results to compute expected negative control values for a range of often-used metrics (Section 4). We furthermore propose an exact test of overall skeleton fit (Section 5), and provide an example of its use on real data. Secondly, we consider more general metrics that are not only concerned with skeleton estimation, and propose a negative control pipeline for this case (Section 6). We provide two examples of its use, both in a simulation study and in a real data application.
@@
But before we turn to these general results, we present an example case where well-known metrics such as adjacency precision and recall do perhaps not behave exactly as one would have expected.
Code for all computations is available online at https://github.com/annennenne/negcontrol-disco.
@@ 2 Precision and Recall: A Cautionary Tale
Table 1: Adjacency Confusion Matrix for the 5 Node DAG Example in Figure 1.
@@
Estimate Adjacency $\textit{tp}=6$ $\textit{fp}=1$
Non-adjacency $\textit{fn}=2$ $\textit{tn}=1$
Consider the two DAGs in Figure 1. The left graph (a) is the true DAG, and the right graph (b) is an estimate produced by a causal discovery procedure. We compute their adjacency confusion matrix in order to evaluate the performance of the discovery procedure (Table 1). This results in:
precision $=\frac{\textit{tp}}{\textit{tp}+\textit{fp}}=\frac{6}{7}\simeq 0.86\quad\quad$ and
recall $=\frac{\textit{tp}}{\textit{tp}+\textit{fn}}=\frac{6}{8}=0.75.$
Are these numbers high or low? Although these values are not too far off from the performance of well-established causal discovery algorithms on simulated data (and much better than typical performance on "real" benchmarking datasets), we will argue that they are indeed as low as can be for this specific discovery task — because the "discovery algorithm" applied here was simply random guessing and hence had absolutely no information about the true data generating mechanism.
More specifically, we simulated 1000 random Erdős-Rényi type DAGs over 5 nodes each with 7 edges [^1] Note that this is not even the correct number of edges, although close to it, as the true DAG has 8 edges. and used these "random guesses" as estimates of the DAG in Figure 1 (a). This resulted in a median precision of 0.86 and a median recall of 0.75, i.e., numbers that exactly match the performance of the example just described. The DAG shown in Figure 1 (b) was one among many random draws that matches this median performance. Hence the large values of precision and recall cannot be attributed to a conveniently chosen random seed.
Is it then a curious artefact for very dense graphs? Or "small" graphs over e.g., 5 nodes? Neither is the case. As we will show in the following section, the phenomenon does not depend on the number of nodes, and depending on the choice of metrics, can occur also in modestly dense graphs.
@@ 3 Distributional Results for Adjacency Metrics Under Random Guessing
Table 2: Generically Labelled Adjacency Confusion Matrix.
@@
Est. Adjacency TP FP $m_{\text{est}}$
Non-adjacency FN TN -
@@
[…]
@@
Consider a DAG $G$ over $d$ nodes, and let $m_{\text{true}}$ denote the number of edges in $G$. Let $\hat{G}$ be another DAG over the same $d$ nodes used as an estimate of $G$, and let $m_{\text{est}}$ be the number of edges in $\hat{G}$. Finally, let $m_{\text{max}}=\sum_{i=1}^{d-1}i=\frac{1}{2}(d-1)d$ denote the maximal number of possible edges in a DAG over $d$ nodes (corresponding to a fully connected graph). We can describe the performance of $\hat{G}$ as an adjacency/skeleton estimator of $G$ through a (generic) confusion matrix as seen in Table 2. Note that for a given causal discovery problem, namely estimating some given $G$, $m_{\text{max}}$ and $m_{\text{true}}$ can be considered fixed: $m_{\text{max}}$ depends only on the number of nodes $d$, which does not change, and $m_{\text{true}}$ is fixed given $G$. Moreover, for many causal discovery procedures, it further makes sense to consider $m_{\text{est}}$ fixed — at least for a specific value of a tuning parameter (e.g., significance level for testing or penalty for a score) and a specific dataset — as we most often do not try to estimate the correct number of edges from data. Rather, algorithms are typically applied with a pre-specified value of the tuning parameter (e.g. significance level of $\alpha=0.01$), which indirectly controls the resulting number of edges, $m_{\text{est}}$ (see Sections 6 and 7 for considerations in cases where this latter assumption is not meaningful).
We now make the following important observation: If edges are placed uniformly in both $G$ and $\hat{G}$ (corresponding to Erdős-Rényi type graphs), and we condition on the row and column sums of Table 2 (which is equivalent with conditioning on $m_{\text{true}}$, $m_{\text{est}}$ and $m_{\text{max}}$), then by definition, the number of true positives will follow a hypergeometric distribution parameterized by $m_{\text{max}}$, $m_{\text{true}}$ and $m_{\text{est}}$:
$\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}}).$
Note that this is an exact distributional result, not an asymptotic statement. This distributional result is well known from its use in the (one-sided) Fisher’s exact test, and may also be motivated using a random urn experiment analogy, see Supplementary Materials A.
@@
This observation gives rise to several useful applications: First, we can compute the expected value, median and uncertainty estimates (e.g., confidence interval) for the number of true positive adjacencies under random guessing. Secondly, since we are also conditioning on $m_{\text{max}}$, $m_{\text{true}}$ and $m_{\text{est}}$, we can further compute expectations and draw statistical inference for any function of the confusion matrix, including precision, recall and F1. We provide formulas for these in Section 4. Thirdly, we can construct an exact statistical test of overall skeleton fit by considering how much the number of true positives in a given estimated graph diverts from its expected distribution under a null hypothesis of random edge placement. We propose such a test in Section 5.
@@ 4 Expectations and Quantiles of Adjacency Metrics Under Random Guessing
Table 3: Expected Values and Quantile Expressions Under Random Guessing for Five Commonly Used Adjacency Metrics.
@@
Precision $\frac{m_{\text{true}}}{m_{\text{max}}}$ $\frac{q_{k}}{m_{\text{est}}}$
Recall $\frac{m_{\text{est}}}{m_{\text{max}}}$ $\frac{q_{k}}{m_{\text{true}}}$
F1 $\frac{2\cdot m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}\cdot m_{\text{est}}+m_{\text{max}}\cdot m_{\text{true}}}$ $\frac{2\cdot q_{k}}{m_{\text{est}}+m_{\text{true}}}$
NPV $1-\frac{m_{\text{true}}}{m_{\text{max}}}$ $\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{k}}{m_{\text{max}}-m_{\text{est}}}$
Specificity $1-\frac{m_{\text{est}}}{m_{\text{max}}}$ $\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{k}}{m_{\text{max}}-m_{\text{true}}}$
[…]
Since
$\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}}),$
by definition we have that
$\text{E}(\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}$
and by considering the quantile function of $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, we can construct a confidence interval as e.g., the central 95% of the distribution, or find the expected median.
For fixed values of $(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, Table 3 provides an overview of expected values and quantiles under random guessing for five metrics commonly used for evaluating adjacency placement for causal discovery algorithms, namely precision, recall, F1 score, negative predictive value (NPV) and specificity. As an example, we here showcase derivations for precision, and refer to Supplementary Materials B for derivations for the remaining four metrics.
@@
We first express precision as a function of $\textit{TP},m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
prec $=\frac{\textit{TP}}{\textit{TP}+\textit{FP}}=\frac{\textit{TP}}{\textit{TP}+m_{\text{est}}-\textit{TP}}=\frac{\textit{TP}}{m_{\text{est}}}$
@@
$\text{E}(\text{prec}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{1}{m_{\text{est}}}\text{E}(\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
@@
The linearity also makes it easy to obtain e.g., an exact 95% confidence interval under the null hypothesis of random guessing by simply applying the same transformation to the appropriate quantiles of $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$. For example, an exact 95% confidence interval for precision under the null is given by
$\left(\frac{1}{m_{\text{est}}}q_{(0.025,m_{\text{max}},m_{\text{true}},m_{\text{est}})},\frac{1}{m_{\text{est}}}q_{(0.975,m_{\text{max}},m_{\text{true}},m_{\text{est}})}\right)$
@@
$\text{median}(\text{prec})=\frac{1}{m_{\text{est}}}q_{(0.5,m_{\text{max}},m_{\text{true}},m_{\text{est}})}.$
General remarks concerning Table 3
@@
Moreover, we note that under random guessing, the expected precision does not depend on the number of edges in the estimated graph ($m_{\text{est}}$), only on the number of edges in the true graph ($m_{\text{true}}$) and the maximal possible number of edges ($m_{\text{max}}$). But recall increases linearly as a function of the number of estimated edges. Hence, if we are using random guessing, a "free lunch" in optimizing precision and recall is achievable simply by estimating a very large number of edges, even including the trivial fully connected graph. This can also be seen from the expected value of the F1 score under random guessing, which increases monotonically with the number of estimated edges: It is always better to just add another edge.
Note that all statistical inference based on Table 3 relies on a single distributional result. Hence, while we may construct confidence intervals for e.g., both precision and recall to aid our own interpretation of a specific estimation problem, there is no additional statistical information gained: We might as well just draw inference directly on the number of true positives.
@@
4.1 Example: Expected Adjacency Precision and Recall for a Dense 5 Node DAG